OpenStudy (anonymous):

$6500 is invested in an account that pays 8 % compounded every 4 months. What is the amount, in the account after 10 years?

OpenStudy (anonymous):

Compound Interest P is the principal (the money you start with, your first deposit) r is the annual rate of interest as a decimal (5% means r = 0.05) n is the number of years you leave it on deposit A is how much money you've accumulated after n years, including interest. If the interest is compounded once a year: A = P(1 + r)n If the interest is compounded q times a year: A = P(1 + r/q)nq Use the second equation: Therefore A = 6500(1+0.08/3)^13 = $9158.58

OpenStudy (anonymous):

It's unclear whether that 8% is per year or per 4 months. I'll assume it's per year. \[A=P(1+r)^{t}\]where P is the principal, r is the rate (over the compounding period), and t is the number of compounding periods. P=$6500 r = 8%/3 t = (10 years)(3 periods/year) = 30 periods \[$6500(1+\frac{0.08}{3})^{30}\]